OpenStudy (anonymous):

Which is the 23rd term of the arithmetic series 5, 17, 29, ... ? 181 203 247 269

OpenStudy (anonymous):

So a1 = 5 d is 12 (it increases by 12 every term) so an = a1 + d (n-1) a23 = 5 + 12(22) a23 = 5+ 264 a23 = 269 that's one of the answers.